OpenStudy (anonymous):

Solve using substitution: 2x = 1 - y 5x - 10y = -10 I NEED HELP

OpenStudy (lalaly):

using the first equation y=1-2x ...(1) first substitute (1) in 5x-10y=-10 to find x can you do that?

OpenStudy (anonymous):

Can you go through this one step by step so I can use it to help with others??

OpenStudy (lalaly):

okay thats what im doing :D

OpenStudy (lalaly):

solving equations using substitution first step you choose one of the equations u have and solve it for x or for y .. but you always choose the easier one,

OpenStudy (lalaly):

in this case, the easiest is the first equation , and since the coeffecient of y is 1 so solve for y

OpenStudy (lalaly):

2x=1-y becomes y=1-2x

OpenStudy (lalaly):

are you following?

OpenStudy (anonymous):

Yes kind of :)

OpenStudy (lalaly):

okay the next step you substitute y=1-2x in y in the second equation which is 5x-10y=-10 substituting it the equation becomes 5x-10(1-2x)=-10

OpenStudy (lalaly):

this way you find x

OpenStudy (lalaly):

solve for x 5x-10+20x=-10 25x=0 x=0 then subtsitute x=0 in any one of the equations to find y

OpenStudy (lalaly):

for example using 2x=1-y 2(0)=1-y 0=1-y y=-1

OpenStudy (anonymous):

Ok thank you!!:)